# Displacement techniques in bounded cohomology

### v1

@@ 1. Introduction
The study of the vanishing of bounded cohomology $H_{b}^{\bullet}(-;-)$ has attracted a lot of interest from the very beginning of the theory, since it is a useful tool in geometric topology [Gro82] and rigidity theory [BM02], and it characterizes certain algebraic properties of groups. The first result in this direction is the celebrated characterization of amenable groups by Johnson [Joh72]: A group $\Gamma$ is amenable if and only if its bounded cohomology vanishes for all dual Banach $\Gamma$-modules and all positive degrees. Similarly, one can characterize finite groups as the ones whose bounded cohomology vanishes in all positive degrees and for all Banach modules [Fri17]. The vanishing of the bounded cohomology in low degrees and trivial real coefficients (i.e. endowed with the trivial action) has already strong algebraic consequences. For instance, via Bavard Duality [Bav91] the stable commutator length of a group $\Gamma$ is zero if $H_{b}^{2}(\Gamma;\mathbb{R})=0$. Finally, groups with trivial bounded cohomology in all positive degrees can be used to compute explicitly some full bounded cohomology rings [FFLM23, Mon22, CFFLM23]. This is the case for example of Thompson’s group $F$, whose bounded cohomology is trivial in positive degrees and real coefficients, that can be used to compute the bounded cohomology ring of Thompon’s group $T$: $H_{b}^{\bullet}(T;\mathbb{R})=\mathbb{R}[x]$, where $x$ denotes the Euler class [FFLM23, Mon22].
@@
Many vanishing results for the (bounded) cohomology available in the literature are based on displacement techniques. Such approaches for ordinary (co)homology date back to works by Anderson [And61] and Fisher [Fis60] in the context of homeomorphism groups, as well as to Mather’s proof of the acyclicity of the compactly supported homeomorphisms group of $\mathbb{R}^{n}$, $H_{k\geq 1}(\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n});\mathbb{Z})=0$ [Mat71]; and to Wagoner’s work on classifying spaces in algebraic $K$-theory  [Wag72]. Based on the work of Mather, Baumslag–Dyer–Heller introduced a large class of acyclic groups called mitotic groups [BDH80] and then Varadarajan [Var85] and Berrick [Ber89] independently extended that class to the larger class of binate groups (that includes dissipated groups such as $\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n})$). In the bounded context the first displacement techniques appeared in the work by Matsumoto and Morita about the bounded acyclicity, i.e. having vanishing bounded cohomology for all positive degrees and trivial real coefficients, of the group $\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n})$ [MM85]. Later Kotschick introduced the notion of commuting conjugates [Kot08, FFL23] a displacement property of a group $\Gamma$ that readily implies the vanishing of the stable commutator length of $\Gamma$. Kotschick’s result has been recently extended by showing that the property of having commuting conjugates actually implies the vanishing of the second bounded cohomology group with trivial real coefficients [FFL23]. Moreover, approaches close to the one of Matsumoto and Morita also show that binate groups are boundedly acyclic [Löh17, FFLM23]. Finally, some new displacement techniques have been recently introduced in the study of $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups, i.e. groups with zero bounded cohomology in all positive degrees and all separable dual modules [Mon22, CFFLM23]. These new dynamical properties are called commuting cyclic conjugates, or commuting $\mathbb{Z}$-conjugates in some special cases [CFFLM23]. Moreover, for convenience we also give a name here to the property previously introduced by Monod [Mon22] of having conjugates in commuting $\mathbb{Z}$-conjugate.
@@ Acknowledgements
This work has been supported by the Madrid Government (Comunidad de Madrid - Spain) under the multiannual Agreement with UAM in the line for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). This work has been funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, from title “Geometry and topolgy of manifolds”, proposal code 2022NMPLT8 - CUP J53D23003820001.
@@ 2. Definitions and inclusions
In this note, $\Gamma$ will always denote an abstract discrete group.
@@ 2.1. Bounded cohomology and stable commutator length
Given a normed $\Gamma$-module $E$, the bounded cohomology of $\Gamma$ with coefficients in $E$, denoted by $H_{b}^{\bullet}(\Gamma;E)$, is defined as the cohomology of the bounded cochain complex $(\ell^{\infty}(\Gamma^{\bullet+1};E)^{\Gamma};\delta^{\bullet})$, where $\delta^{\bullet}$ denotes the standard simplicial coboundary operator (which sends bounded functions to bounded functions).
A classical result by Johnson shows that amenable groups can be characterised as those groups with trivial bounded cohomology in all positive degrees and all dual coefficients [Joh72]. On the other hand, the vanishing of the degree two bounded cohomology with trivial real coefficients is much more common than just amenable groups, and provides useful information about the algebraic structure of the group. Indeed, given a group $\Gamma$ we can define the commutator length of an element $g\in\Gamma^{\prime}\coloneqq[\Gamma,\Gamma]\leq\Gamma$, denoted by $\textup{cl}(g)$, as the minimal integer $n\in\mathbb{N}$ such that $g$ can be written as the product of $n$ commutators. The stable commutator length of an element $g\in\Gamma^{\prime}$ is then defined as follows [Cal09]:
@@
Bavard [Bav91] showed that the vanishing of the stable commutator length on $\Gamma^{\prime}$ is equivalent to the injectivity of the comparison map in degree two:
@@
which is the map induced by the inclusion of the bounded cochain complex into the ordinary one. This result witnesses that groups with trivial second bounded cohomology with real coefficients have vanishing stable commutator length.
The vanishing of second bounded cohomology with trivial real coefficients is more robust than the vanishing of stable commutator length. For instance it is stable under taking central extensions, something that is of crucial importance in the study of the spectrum of simplicial volume [HL21, HL19, FFL23]. Moreover, the vanishing of second bounded cohomology with trivial real coefficients has strong consequences for the rigidity of circle actions [Ghy01, Mat86, Bur11].
@@ 2.2. Boundedly acyclic groups
In this note we are interested in boundedly acyclic groups [MR23, CFFLM23], i.e. groups with vanishing bounded cohomology in positive degrees:
@@
A group $\Gamma$ is said to be boundedly $n$-acyclic if $H_{b}^{k}(\Gamma;\mathbb{R})=0$ for all integers $1\leq k\leq n$. We say that $\Gamma$ is boundedly acyclic if $n=\infty$.
Similarly, a group $\Gamma$ is said to be $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic if $H^{k}_{b}(\Gamma;E)=0$ for all integers $k>0$ and all separable dual Banach $\Gamma$-modules $E$.
@@ 2.2.1. Binate, mitotic and dissipated groups
We say that $\Gamma$ is binate if for every finitely generated subgroup $H\leq\Gamma$, there exists a homomorphism $f\colon H\to\Gamma$ such that $[H,f(H)]=1$, and an element $t\in\Gamma$ such that ${}^{t}f(h)=h\cdot f(h)$ for all $h\in H$.
@@
Since binate groups are acyclic [Ber89], it is sufficient to show that the comparison map $\operatorname{comp}^{k}\colon H^{k}_{b}(\Gamma;\mathbb{R})\to H^{k}(\Gamma;\mathbb{R})$ is injective in all positive degrees in order to conclude that binate groups are boundedly acyclic. This is the strategy initiated by Matsumoto and Morita [MM85] in the case of compactly supported homeomorphisms of $\mathbb{R}^{n}$ and extended by several authors that led to the following:
Theorem 2.4 (Binate groups are boundedly acyclic [FFLM23]).
Binate groups are boundedly acyclic.
@@
We say that $\Gamma$ is mitotic if for every finitely generated subgroup $H\leq\Gamma$ there exist $t_{1},t_{2}\in\Gamma$ such that $[H,{}^{t_{1}}H]=1$ and ${}^{t_{2}}h=h\cdot{}^{t_{1}}h$ for all $h\in H$.
In other words, the homomorphism $f\colon H\to\Gamma$ in the definition of binate groups is required to be a conjugate of the given embedding. The bounded acyclicity of such groups was already proved by Löh [Löh17].
Mitotic groups are very suitable for constructions in combinatorial group theory. However, they do not appear naturally in geometry and dynamics (we will discuss this more in Section 3.5). A much more common class of binate groups which contains the most natural examples of homeomorphism groups, is the class of dissipated groups:
@@
Let $\Gamma$ be a group of bijections of a set $X$, written as a directed union of a direct system of subsets $(X_{i})_{i\in I}$, which we call bounded sets. We say that $\Gamma$ is boundedly supported if every element $g\in\Gamma$ satisfies $\operatorname{supp}(g)\subset X_{i}$ for some $i\in I$.
@@
If there exists a dissipator for $\Gamma$ and $X_{i}$ for every $i\in I$, we say that $\Gamma$ is dissipated.
@@
Dissipated groups typically arise as groups of compactly supported homeomorphisms [SV90]. Historically, the most important example is the group of compactly supported homeomorphisms of $\mathbb{R}^{n}$: Indeed, based on the work by Mather in the case of ordinary (co)homology [Mat71], Matsmumoto and Morita proved that the comparison map is injective in all positive degrees [MM85]. This was the first example of a non-amenable boundedly acyclic group.
@@ 2.2.2. Commuting conjugates, second bounded cohomology and stable commutator length
The next displacement property that appeared in the literature is the one of commuting conjugates, which is the weakest and most natural one. This notion was introduced by Kotschick [Kot08, Proposition 2.2] for studying the stable commutator length, and then deeply investigated in the context of second bounded cohomology [FFL23]. Kotschick’s definition is stated in a slightly different way but clearly equivalent to the following:
@@
We say that $\Gamma$ has commuting conjugates (abbreviated c. c.) if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t}H]=1$.
@@
Let $\Gamma$ be a group with commuting conjugates. Then the stable commutator length vanishes identically on Γsuperscriptnormal-$\Gamma^{\prime}$.
This was then strengthened to the vanishing of the whole second bounded cohomology group:
@@
We now move to the properties in Figure 1. For a group to have commuting conjugates, there needs to be no extra structure on the sets of elements $t$ that appear in the definition. However, in order to obtain the vanishing of bounded cohomology in higher degrees, we have to require that the sets of conjugates are well organized. This leads to the definition of commuting cyclic conjugates:
@@
We say that $\Gamma$ has commuting cyclic conjugates (abbreviated c. c. c.) if for every finitely generated subgroup $H\leq\Gamma$ there exist $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $[H,t^{n}]=1$. Here we read that $t^{\infty}=1$.
@@
We say that $\Gamma$ has commuting $\mathbb{Z}$-conjugates (abbreviated c. $\mathbb{Z}$ c.) if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t^{p}}H]=1$ for all $p\geq 1$.
@@
The main advantage of commuting $\mathbb{Z}$-conjugates over dissipated is that in the former case the existence of elements acting simultaneously on infinitely many components of support is not required. As described by Monod, this recalls the Aristotelian distinction between actual and potential infinity [Mon22, p. 664]. As a specific instance of this, certain groups of diffeomorphisms or piecewise linear homeomorphisms have commuting $\mathbb{Z}$-conjugates, but they cannot be dissipated, because of the singularity that forms at the accumulation point of the shifted copies of bounded sets [Kot08, p. 402].
@@
Let $n\geq 2$. We say that $\Gamma$ has commuting $\mathbb{Z}/n$-conjugates (abbreviated c. $\mathbb{Z}/n$ c.) if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$, and $[H,t^{n}]=1$.
@@
As anticipated before, the notion of commuting cyclic conjugates, which slightly strengthens the notion of commuting conjugates, implies a much stronger vanishing result:

### v2

@@ 1. Introduction
The study of the vanishing of bounded cohomology $H_{b}^{\bullet}(-;-)$ has attracted a lot of interest from the very beginning of the theory, since it is a useful tool in geometric topology [Gro82] and rigidity theory [BM02], and it characterises certain algebraic properties of groups. The first result in this direction is the celebrated characterization of amenable groups by Johnson [Joh72]: A group $\Gamma$ is amenable if and only if its bounded cohomology is zero for all dual Banach $\Gamma$-modules and all positive degrees. Similarly, one can characterise finite groups as the ones whose bounded cohomology is zero in all positive degrees and for all Banach modules [Fri17]. The vanishing of the bounded cohomology in low degrees and trivial real coefficients (i.e. endowed with the trivial action) has already strong algebraic consequences. For instance, via Bavard Duality [Bav91] the stable commutator length of a group $\Gamma$ is zero if $H_{b}^{2}(\Gamma;\mathbb{R})=0$. Finally, groups with trivial bounded cohomology in all positive degrees can be used to compute explicitly some full bounded cohomology rings [Mon22, FFLM23, CFFLM23]. This is the case for example of Thompson’s group $F$, whose bounded cohomology is trivial in positive degrees and real coefficients. This fact can be used to compute the bounded cohomology ring of Thompson’s group $T$, leading to $H_{b}^{\bullet}(T;\mathbb{R})=\mathbb{R}[x]$, where $x$ denotes the Euler class [Mon22, FFLM23].
@@
Many vanishing results for the (bounded) cohomology available in the literature are based on displacement techniques. Such approaches for ordinary (co)homology date back to works by Anderson [And61] and Fisher [Fis60] in the context of homeomorphism groups, as well as to Mather’s proof of the acyclicity of the compactly supported homeomorphisms group of $\mathbb{R}^{n}$, $H_{k\geq 1}(\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n});\mathbb{Z})=0$ [Mat71]; and to Wagoner’s work on classifying spaces in algebraic $K$-theory  [Wag72]. Based on the work of Mather, Baumslag–Dyer–Heller introduced a large class of acyclic groups called mitotic groups [BDH80]. Then Varadarajan [Var85] and Berrick [Ber89] independently extended that class to the larger family of binate groups (that includes dissipated groups such as $\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n})$). In the bounded context the first displacement techniques appeared in the work by Matsumoto and Morita about the bounded acyclicity of the group $\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{n})$ [MM85] (recall that a group is boundedly acyclic if its bounded cohomology is zero in all positive degrees and trivial real coefficients). Later Kotschick introduced the notion of commuting conjugates [Kot08, FFL23], a displacement property of a group $\Gamma$ that readily implies the vanishing of the stable commutator length of $\Gamma$. Kotschick’s result has been recently extended by showing that the property of having commuting conjugates actually implies that the second bounded cohomology group with trivial real coefficients is zero [FFL23]. Moreover, approaches close to the one of Matsumoto and Morita also show that binate groups are boundedly acyclic [Löh17, FFLM23]. Finally, some new displacement techniques have been recently introduced in the study of $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups, i.e. groups with zero bounded cohomology in all positive degrees and all separable dual modules [Mon22, CFFLM23]. These new dynamical properties are called commuting cyclic conjugates, or commuting $\mathbb{Z}$-conjugates in some special cases [CFFLM23]. A similar dynamical property appears in the work of Monod [Mon22], which we refer to as having conjugates in commuting $\mathbb{Z}$-conjugate.
@@ Acknowledgements
This work has been supported by the Madrid Government (Comunidad de Madrid - Spain) under the multiannual Agreement with UAM in the line for the Excellence of the University Research Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). This work has been funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, with title “Geometry and topology of manifolds”, proposal code 2022NMPLT8 - CUP J53D23003820001.
@@ 2. Definitions and inclusions
In this note, $\Gamma$ will always denote an abstract discrete group. Some of the material recalled here is standard, and more detail can be found in e.g. [Mon01, Cal09, Fri17].
@@ 2.1. Bounded cohomology and stable commutator length
Given a normed $\Gamma$-module $E$, the bounded cohomology of $\Gamma$ with coefficients in $E$, denoted by $H_{b}^{\bullet}(\Gamma;E)$, is defined as the cohomology of the bounded cochain complex $(\ell^{\infty}(\Gamma^{\bullet+1};E)^{\Gamma};\delta^{\bullet})$, where $\delta^{\bullet}$ denotes the standard simplicial coboundary operator (which sends bounded functions to bounded functions) and $\ell^{\infty}(\Gamma^{\bullet+1};E)^{\Gamma}$ denotes the subcomplex of invariant functions of $\ell^{\infty}(\Gamma^{\bullet+1};E)$.
A classical result by Johnson shows that amenable groups can be characterised as those with trivial bounded cohomology in all positive degrees and all dual coefficients [Joh72]. On the other hand, the vanishing of the degree two bounded cohomology with trivial real coefficients is much more common than amenability, and provides useful information about the algebraic structure of the group. Indeed, given a group $\Gamma$ we can define the commutator length of an element $g\in\Gamma^{\prime}\coloneqq[\Gamma,\Gamma]\leq\Gamma$, denoted by $\textup{cl}(g)$, as the minimal integer $n\in\mathbb{N}$ such that $g$ can be written as the product of $n$ commutators. The stable commutator length of an element $g\in\Gamma^{\prime}$ is then defined as follows [Cal09]:
@@
Bavard [Bav91] showed that the the stable commutator length is identically zero on $\Gamma^{\prime}$ if and only if the comparison map in degree two:
@@
which is the map induced by the inclusion of the bounded cochain complex into the ordinary one, is injective. This result witnesses that groups with zero second bounded cohomology with real coefficients have zero stable commutator length.
The vanishing of second bounded cohomology with trivial real coefficients is more robust than the vanishing of stable commutator length. For instance, the fact that the vanishing of second bounded cohomology is stable under taking central extensions, while the vanishing of stable commutator length is not, is of crucial importance in the study of the spectrum of simplicial volume [HL21, FFL23, HL24]. Moreover, if the second bounded cohomology with trivial real coefficients is zero, one can deduce strong consequences for the rigidity of circle actions [Mat86, Ghy01, Bur11].
@@ 2.2. Boundedly acyclic groups
In this note we are interested in boundedly acyclic groups [MR23, CFFLM23], i.e. groups with zero bounded cohomology in positive degrees:
@@
A group $\Gamma$ is boundedly $n$-acyclic if $H_{b}^{k}(\Gamma;\mathbb{R})=0$ for all integers $1\leq k\leq n$. It is boundedly acyclic if $n=\infty$.
Similarly, a group $\Gamma$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic if $H^{k}_{b}(\Gamma;E)=0$ for all integers $k>0$ and all separable dual Banach $\Gamma$-modules $E$.
Recall that by Johnson’s Theorem amenable groups are characterised by the vanishing of bounded cohomology with all dual coefficients. Therefore $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity is in some sense the strongest vanishing property that one can hope for in non-amenable groups. Still, it has very strong consequences in rigidity theory, see [CFFLM23, Section 6].
@@ 2.2.1. Binate, mitotic and dissipated groups
A group $\Gamma$ is binate if for every finitely generated subgroup $H\leq\Gamma$, there exists a homomorphism $f\colon H\to\Gamma$ such that $[H,f(H)]=1$, and an element $t\in\Gamma$ such that ${}^{t}f(h)=h\cdot f(h)$ for all $h\in H$.
@@
A group $\Gamma$ is mitotic if for every finitely generated subgroup $H\leq\Gamma$ there exist $t_{1},t_{2}\in\Gamma$ such that $[H,{}^{t_{1}}H]=1$ and ${}^{t_{2}}h=h\cdot{}^{t_{1}}h$ for all $h\in H$.
In other words, the homomorphism $f\colon H\to\Gamma$ in the definition of binate groups is required to be a conjugate of the given embedding.
Mitotic groups are very suitable for constructions in combinatorial group theory. However, they do not appear naturally in geometry and dynamics (we will discuss this more in Section 3.5). A much more common class of binate groups, which contains the most natural examples of homeomorphism groups, is the class of dissipated groups:
@@
Let $\Gamma$ be a group of bijections of a set $X$, written as a directed union of a direct system of subsets $(X_{i})_{i\in I}$, which we call bounded sets. The group $\Gamma$ is boundedly supported if every element $g\in\Gamma$ satisfies $\operatorname{supp}(g)\subset X_{i}$ for some $i\in I$.
@@
The group $\Gamma$ is dissipated if there exists a dissipator for $\Gamma$ and $X_{i}$ for every $i\in I$.
@@
Dissipated groups typically arise as groups of compactly supported homeomorphisms [SV90]. Historically, the most important example in the context of bounded cohomology is the group of compactly supported homeomorphisms of $\mathbb{R}^{n}$: Indeed, based on the work by Mather in the case of ordinary (co)homology [Mat71], Matsmumoto and Morita proved that the comparison map $\operatorname{comp}^{k}\colon H^{k}_{b}(\Gamma;\mathbb{R})\to H^{k}(\Gamma;\mathbb{R})$ is injective in all positive degrees [MM85]. This was the first example of a non-amenable boundedly acyclic group.
The same strategy then became available for binate groups: since they are acyclic [Ber89], it is sufficient to show that their comparison map is injective in all positive degrees in order to conclude that binate groups are boundedly acyclic. This approach was first applied by Löh to the case of mitotic groups [Löh17] and then more recently to all binate groups:
Theorem 2.8 (Binate groups are boundedly acyclic [FFLM23]).
Binate groups are boundedly acyclic.
@@ 2.2.2. Commuting conjugates, second bounded cohomology and stable commutator length
The next displacement property that appeared in the literature is the one of commuting conjugates, which is the weakest and most natural one. This notion was introduced by Kotschick [Kot08, Proposition 2.2] for studying the stable commutator length, and then deeply investigated in the context of second bounded cohomology [FFL23]. Kotschick’s results are phrased with slightly different assumptions [Kot08, Proposition 2.1], but the underlying property is the following:
@@
The group $\Gamma$ has commuting conjugates (abbreviated c. c.) if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t}H]=1$.
@@
Let $\Gamma$ be a group with commuting conjugates. Then the stable commutator length vanishes identically on $\Gamma^{\prime}$.
As we mentioned earlier, by Bavard Duality [Bav91] this result is equivalent to the injectivity of the comparison map in degree two [Fri17, Corollary 2.11]. This was then strengthened to the vanishing of the whole second bounded cohomology group:
@@
We now move to the properties in Figure 1. For a group to have commuting conjugates, there needs to be no extra structure on the sets of elements $t$ that appear in the definition. However, in order to obtain the vanishing of bounded cohomology in higher degrees, we have to require that the sets of conjugates are well organized  [CFFLM23]. This leads to the definition of commuting cyclic conjugates:
@@
The group $\Gamma$ has commuting cyclic conjugates (abbreviated c. c. c.) if for every finitely generated subgroup $H\leq\Gamma$ there exist $t\in\Gamma$ and $n\in\mathbb{Z}_{\geq 2}\cup\{\infty\}$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$ and $[H,t^{n}]=1$. Here we read that $t^{\infty}=1$.
@@
The group $\Gamma$ has commuting $\mathbb{Z}$-conjugates (abbreviated c. $\mathbb{Z}$ c.) if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t^{p}}H]=1$ for all $p\geq 1$.
@@
The main advantage of commuting $\mathbb{Z}$-conjugates over dissipated is that in the former case the existence of elements acting simultaneously on infinitely many components of support is not required. This recalls the Aristotelian distinction between actual and potential infinity [Mon22, p. 664]. As a specific instance of this, certain groups of diffeomorphisms or piecewise linear homeomorphisms have commuting $\mathbb{Z}$-conjugates, but they cannot be dissipated, because of the singularity that forms at the accumulation point of the shifted copies of bounded sets [Kot08, p. 402].
@@
Let $n\geq 2$. The group $\Gamma$ has commuting $\mathbb{Z}/n$-conjugates (abbreviated c. $\mathbb{Z}/n$ c.) if for every finitely generated subgroup $H\leq\Gamma$ there exists $t\in\Gamma$ such that $[H,{}^{t^{p}}H]=1$ for $1\leq p<n$, and $[H,t^{n}]=1$.
@@
Recall that groups with commuting conjugates have zero second bounded cohomology (Theorem 2.12). As anticipated before, the notion of commuting cyclic conjugates, which slightly strengthens the notion of commuting conjugates, implies a much stronger vanishing result:
@@
Note that it is not known whether all binate groups are $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic (Problem 4.1).